OpenStudy (anonymous):

Pastaroni is a local Italian restaurant that serves a specialized pasta salad. Included in this salad is macaroni, linguine and bow-tie pasta. The salad contains six times as much linguine as it does macaroni. Both macaroni and linguine sell for $4; bow-tie pasta sells for $8. If the restaurant makes 140 pounds of pasta every day, how many pounds of each type of pasta are needed if the salad sells for $6?

OpenStudy (anonymous):

Let M, L and BT be the pound weight of macaroni, linfuine and bow-tie pasta respectively. Solve the following equations for M, L and BT:\[\left\{L+M+\text{BT}=140,\frac{4L+4M+8\text{BT}}{140}=6,\frac{L}{M}=6\right\}\]\[\{L=60,M=10,\text{BT}=70\} \]Verify the result:\[\frac{60*\text{$\$$4}+10*\text{$\$$4}+70*\text{$\$$8}}{140}=\text{$\$$6} \]

OpenStudy (anonymous):

Sorry, linfuine should have been written, linguine.

OpenStudy (anonymous):

THANKS!!

OpenStudy (anonymous):

You are welcome and thank you for the medal.

OpenStudy (anonymous):

Okay wait, how did you get L=60 M=10 and b=70

OpenStudy (anonymous):

@CamW Refer to the attached Mathematica calculation to see how the values of L, M and BT were determined.